The expenses for a certain journey are composed of train, bus snd taxi fares. The taxi fare is 10 times the bus fare and the train fare is 4times the bus fare. The total expenditure is $30. Find the taxi fare. The taxi fare is 'b'. Show your steps.
Answer by mananth(16946)

Taxi fare = b
bus fare = b/10
train fare = 4b/10
b+ b/10 + 4b/10 = 30
multiply by LCD = 30
30b+3b+12b=900
45b=900
b=900/45
b=20
Taxi fare = $20
bus fare = 20/10 = $2
Train fare = 80/10 = $8